# A Puzzle About General Covariance and Gauge

### v2

@@ 7 Solder Forms
The argument is much like before. Since $W^{1}(W^{r-1}P)$ carries the canonical form $\theta_{W^{r-1}G}$, and every local smooth map $f:\mathbb{R}^{n}\times G\rightarrow P$ lifts to a local smooth map $j^{r}f:\mathbb{R}^{n}\times W^{r}(n,\mathbb{R},G)\rightarrow W^{r}P$ in the neighbourhood of the identity element $e^{r}_{G}$ of $W^{r}(n,\mathbb{R},G)$, we can pull $\theta_{W^{r-1}G}$ back to $W^{r}P$ via the map $W^{r}P\rightarrow W^{1}(W^{r-1}P)$, defined by $j^{r}_{(0,e)}f\mapsto j^{1}_{(0,e)}j^{r-1}f$ to obtain $\theta^{r}_{G}$. So again, putting Theorem 2 and Fact 2 together, solder forms are also a generic feature of gauge natural bundles, in the sense that any gauge natural bundle (excepting the 00th-order case) is associated to a bundle which carries a solder form.
@@ 8 Conclusion: General Covariance Revisited
Our primary goal in this paper was to extract, from the long and vexed literature on general covariance, certain precise necessary conditions; and then to argue that those conditions are non-trivial because in fact they are violated by Yang-Mills theory. We then argued that something like general covariance, properly generalized, could be restored for Yang-Mills theory, but only by moving to a more general mathematical setting; and we showed how this perspective could provide fruitful insight into other different debates, concerning the status and significance of principal bundle and solder forms in understanding the relationship between general relativity and Yang-Mills theory.
We have not attempted to give a general or complete account of general covariance. As we have noted, to do so, we would need to say much more about what it means for a theory to have the right sort of behavior under diffeomorphisms, whereas our focus has been on properties of the objects that our theories posit. We postpone that more complete discussion for future work. Even so, we will conclude by suggesting that the (somewhat preliminary) arguments and observations offered here offer a valuable perspective on what general covariance and related issues, such as coordinate-independence, are really about.
@@
We can restore a minimal sense of general covariance for Yang-Mills theory by recognizing that the background geometry of Yang-Mills theory consists of only the structure a principal bundle, along with a family of vector bundles, all systematically related to one another, endowed with certain further structure preserved under the action of the principal bundle’s structure group: an inner product and, in some cases, an orientation. No further structure, such as a preferred trivialization or coordinatization, is assumed. In this, general covariance for Yang-Mills theory is analogous to general covariance for general relativity, where the background geometry consists of only the structure of a smooth manifold, along with various bundles that can be constructed from that. Of course, to represent concrete situations in Yang-Mills theory, one adds to this background structure sections of certain gauge natural bundles, such as a connection and vector fields representing matter configurations. But this is also analogous to general relativity, where one introduces a metric and various matter fields, all sections of natural bundles.
Talk of “structures” and “dependence” is evocative, but not very precise. But the key idea in the present case is explicated using the formalism of natural and gauge-natural bundles. We suggest that it is naturality, or more precisely, still, functoriality, that captures the core of what general covariance is concerned with. The difference between the sorts of objects and relationships we encounter in general relativity – the generally covariant ones – and those in Yang-Mills theory is that the former, but not the latter, involve objects that depend only on a manifolds that represents space and time; whereas the latter depends on further structure on top of that. Functoriality reflects this idea by enforcing the requirement that our constructions can be applied uniformly across manifolds, in a way that is compatible with those maps that we take to preserve manifold structure. Coordinate- or frame-dependent constructions fail to be functorial (over manifolds) because they require further information that is not generally preserved by diffeomorphisms. This is why those constructions fail to be generally covariant. But that is not the only way to fail to be generally covariant, and the natural bundle formalisms shows very clearly why.

### v3

@@ 7 Solder Forms
The argument is much like before. We have seen, by the argument above, that $W^{1}P$ carries a canonical form. Then, for $r\geq 1$, since $W^{1}(W^{r-1}P)$ carries the canonical form $\theta_{W^{r-1}G}$, and every local smooth map $f:\mathbb{R}^{n}\times G\rightarrow P$ lifts to a local smooth map $j^{r}f:\mathbb{R}^{n}\times W^{r}(n,\mathbb{R},G)\rightarrow W^{r}P$ in the neighbourhood of the identity element $e^{r}_{G}$ of $W^{r}(n,\mathbb{R},G)$, we can pull $\theta_{W^{r-1}G}$ back to $W^{r}P$ via the map $W^{r}P\rightarrow W^{1}(W^{r-1}P)$, defined by $j^{r}_{(0,e)}f\mapsto j^{1}_{(0,e)}j^{r-1}f$ to obtain $\theta^{r}_{G}$. So again, putting Theorem 2 and Fact 2 together, solder forms are also a generic feature of gauge natural bundles, in the sense that any gauge natural bundle is associated to a bundle which carries a solder form.
These observation do not quite complete the argument, however, and a further clarification is called for. We have shown that gauge natural bundles of order $r\geq 1$ carry a canonical form. But what about bundles of order 00? [^27] Recall the clarification about order counting in note 26. The tangent bundle is order 1 as a natural bundle, but order 0 as a gauge natural bundle, and so the arguments we give presently apply to that as well. Those are the bundles we encounter most often in Yang-Mills theories, as associated vector bundles to principal bundles. The construction we have given here does not apply to those. Instead, those gain a solder form by the argument given by Weatherall (2016), once we identify the principal bundle to which they are associated as a subbundle of their frame bundle. This is possible, and can be done canonically via the definitions of the structures in question, for any associated vector bundle determined by a faithful representation. The point of the argument above is to show that once one has such a solder form for a 0th order bundle, solder forms exist for all higher order jet bundles over that vector bundle, in a way exactly analogous to how higher order natural bundles comes equipped with solder forms.
@@ 8 Conclusion: General Covariance Revisited
Our primary goal in this paper was to extract, from the long and vexed literature on general covariance, certain precise necessary conditions; and then to argue that those conditions are non-trivial because in fact they are violated by Yang-Mills theory. We then argued that something like general covariance, properly generalized, could be restored for Yang-Mills theory, but only by moving to a more general mathematical setting; and we showed how this perspective could provide fruitful insight into other debates, concerning the status and significance of principal bundle and solder forms in understanding the relationship between general relativity and Yang-Mills theory.
We have not attempted to give a general or complete account of general covariance. As we have noted, to do so, we would need to say much more about what it means for a theory to have the right sort of behavior under diffeomorphisms, whereas our focus has been on properties of the objects that appear in our theories. We postpone that more complete discussion for future work. Even so, we will conclude by suggesting that the (somewhat preliminary) arguments and observations here offer a valuable perspective on what general covariance and related issues, such as coordinate-independence, are really about.
@@
We can restore a minimal sense of general covariance for Yang-Mills theory by recognizing that the background geometry of Yang-Mills theory consists of only the structure of a principal bundle, along with a family of vector bundles, all systematically related to one another, endowed with certain further structure preserved under the action of the principal bundle’s structure group: an inner product and, in some cases, an orientation. No further structure, such as a preferred trivialization or coordinatization, is assumed. In this way, general covariance for Yang-Mills theory is analogous to general covariance for general relativity, where the background geometry consists of only the structure of a smooth manifold, along with various bundles that can be constructed from that. Of course, to represent concrete situations in Yang-Mills theory, one adds to this background structure sections of certain gauge natural bundles, such as a connection and vector fields representing matter configurations. But this is also analogous to general relativity, where one introduces a metric and various matter fields, all sections of natural bundles.
Talk of “structures” and “dependence” is evocative, but not very precise. The key idea in the present case is explicated using the formalism of natural and gauge-natural bundles. We suggest that it is naturality, or more precisely, still, functoriality, that captures the core of what general covariance is concerned with. The difference between the sorts of objects and relationships we encounter in general relativity – the generally covariant ones – and those in Yang-Mills theory is that the former, but not the latter, involve objects that depend only on a manifold that represents space and time; whereas the latter depends on further structure on top of that. Functoriality reflects this idea by enforcing the requirement that our constructions can be applied uniformly across manifolds, in a way that is compatible with those maps that we take to preserve manifold structure. Coordinate- or frame-dependent constructions fail to be functorial (over manifolds) because they require further information that is not generally preserved by diffeomorphisms. This is why those constructions fail to be generally covariant. But that is not the only way to fail to be generally covariant, and the natural bundle formalisms shows very clearly why.